Suppose market demand is given Qd=a-bp as and market supply is given as, Qs=c+dp and then find the following



o Equilibrium price



o Equilibrium quantity



o Price elasticity of demand at equilibrium



o Price elasticity of supply at equilibrium

1
Expert's answer
2021-11-23T10:57:06-0500

a) Equilibrium price;

At equilibrium; "Q_{d}=Q_{s}"

"a-bp=c+dp"

"a-c=p(b+d)"

"P_{e}=\frac{a-c}{b+d}"


b)Equilibrium quantity;

"Q_{e}=a-b(\frac{a-c}{b+d})"

"=\frac{ad+ac}{b+d}"


c) Price elasticity of demand at equilibrium; we will use point elasticity of demand formula

Point Elasticity of demand; ="(\frac{-1}{slope})(\frac{P_{e}}{Q_{e}})"

"\frac{P_{e}}{Q_{e}}=\frac{a-c}{b+d}\cdot\frac{b+d}{ad+ac}=\frac{a-c}{ad+ac}"

The demand equation will be rewritten as;

"Q_{d}=a-bp"

"bp=a-Q_{d}"

"p=\frac{a}{b}-(\frac{1}{b})Q_{d}"

Price elasticity of demand"=(-1.-b)\cdot\frac{a-c}{ad+ac}=\frac{ab-ac}{ad+ac}"


d) Price elasticity of supply at equilibrium; we will use point elasticity of supply formula

Point Elasticity of demand; ="(\frac{-1}{slope})(\frac{P_{e}}{Q_{e}})"

"\frac{P_{e}}{Q_{e}}=\frac{a-c}{b+d}\cdot\frac{b+d}{ad+ac}=\frac{a-c}{ad+ac}"

The supply equation will be rewritten as;

"Q_{d}=c+dp"

"dp=Q_{s}-c"

"p=(\frac{1}{d})Q_{s}-\frac{c}{d}"

Price elasticity of supply"=(1.d)\cdot\frac{a-c}{ad+ac}=\frac{ad-cd}{ad+ac}"